Find the coefficient of quartile deviation from the following:



Class 5 – 10 10 – 15 15 – 20 20 – 25 25 – 30 30 – 35 35 – 40



Frequency 6 8 17 21 15 11 2



9. Find the semi-inter-quartile range for the following data:



Marks 0 – 10 10 – 20 20 – 30 30 – 40 40 – 50



No. of students 22 38 46 35 19



10. Find which of the following industry has less variation if the salaries paid to the employees



are as given below:



Salary (Rs. in



1000’s)



Below



4500



4500 –



4700



4700 –



4900



47900 –



5100



5100 &



above



Industry A 6 13 20 7 6



Industry B 4 9 16 7 0

1
Expert's answer
2021-12-31T06:37:57-0500

8.

"n=80"

Q1 class :

Class with (n/4)th value of the observation in cf (cumulative frequency) column

=(80/4)th value of the observation in cf column

=(20)th value of the observation in cf column

and it lies in the class 15-20.

Q1 class : 15-20

The lower boundary point of 15-20 is 15.

L=15

"Q_1=L+\frac{n/4-cf}{f}c=15+\frac{20-14}{17}\cdot5=16.76"


Q3 class :

Class with (3n/4)th value of the observation in cf (cumulative frequency) column

=(3"\cdot" 80/4)th value of the observation in cf column

=(60)th value of the observation in cf column

and it lies in the class 25-30.

Q3 class : 25-30

The lower boundary point of 25-30 is 25.

L=25


"Q_3=L+\frac{3n/4-cf}{f}c=25+\frac{60-52}{15}\cdot5=27.67"


Coefficient of Quartile deviation:


"\frac{Q_3-Q_1}{Q_3+Q_1}=\frac{27.67-16.76}{27.67+16.76}=0.245"


9.

"n=160"

Q1 class :

Class with (n/4)th value of the observation in cf (cumulative frequency) column

=(160/4)th value of the observation in cf column

=(40)th value of the observation in cf column

and it lies in the class 10-20.

Q1 class : 10-20

The lower boundary point of 10-20 is 10.

L=10

"Q_1=L+\frac{n/4-cf}{f}c=10+\frac{40-22}{38}\cdot10=14.74"


Q3 class :

Class with (3n/4)th value of the observation in cf (cumulative frequency) column

=(3"\cdot" 160/4)th value of the observation in cf column

=(120)th value of the observation in cf column

and it lies in the class 25-30.

Q3 class : 30-40

The lower boundary point of 30-40 is 30.

L=30


"Q_3=L+\frac{3n/4-cf}{f}c=30+\frac{120-106}{35}\cdot10=34"


semi-inter-quartile range:


"\frac{Q_3-Q_1}{2}=\frac{34-14.74}{2}=9.63"


10.

for A:

Mean:

"\mu=\frac{\sum f_ix_i}{n}=4776.92"

Standard deviation:

"\sigma=\sqrt{\frac{\sum f_ix_i^2-(\sum f_ix_i)^2/n}{n}}=227.54"

Coefficient of Variation:

"k=\sigma/\mu=227.54/4776.92=4.76\%"


for B:

Mean:

"\mu=\frac{\sum f_ix_i}{n}=4744.44"

Standard deviation:

"\sigma=\sqrt{\frac{\sum f_ix_i^2-(\sum f_ix_i)^2/n}{n}}=180.19"

Coefficient of Variation:

"k=\sigma/\mu=180.19/4744.44=3.8\%"


so, industry B has less variation